Twenty raffle tickets numbered 1 to 20 and one is drawn at random What is the probability that the number shown on the ticket is an odd number or a prime number?

a. 0.4
b. 0.55
c. 0.9
d. 0.5
1
Expert's answer
2020-12-24T17:15:37-0500

n(S) = {1, 2, 3, 4, 5, 6, 7, 8, 9, 10

11, 12, 13, 14, 15, 16, 17, 18, 19, 20}

Number of odd tickets:

1, 3, 5, 7, 9, 11, 13, 15, 17, 19

n(E1) = 10

Number of prime number tickets:

2, 3, 5, 7, 11, 13, 17, 19

n(E2) = 8

Total event = n(E1) + n(E2) – common element number

"= 10 + 8 - 7"

n(E) = 11

"P = \frac{n(E)}{n(S)} \\

=\frac{11}{20} \\

= 0.55"

Answer: b. 0.55